How to represent 49 in Roman?
Found 2 solutions by Fombitz, CoolRekt:
Answer by Fombitz(32388)

49=40+9=(XL)(IX)=XLIX

Answer by CoolRekt(53)

XLIX = 49
XL = 40
IX = 9
XL + IX = 40 + 9 = XLIX = 49
CoolRekt